# Dark Experience for Incremental Keyword Spotting

### v1

@@ Abstract
Spoken keyword spotting (KWS) is crucial for identifying keywords within audio inputs and is widely used in applications like Apple Siri and Google Home, particularly on edge devices. Current deep learning-based KWS systems, which are typically trained on a limited set of keywords, can suffer from performance degradation when encountering new domains, a challenge often addressed through few-shot fine-tuning. However, this adaptation frequently leads to catastrophic forgetting, where the model’s performance on original data deteriorates. Progressive continual learning (CL) strategies have been proposed to overcome this, but they face limitations such as the need for task-ID information and increased storage, making them less practical for lightweight devices. To address these challenges, we introduce Dark Experience for Keyword Spotting (DE-KWS), a novel CL approach that leverages dark knowledge to distill past experiences throughout the training process. DE-KWS combines rehearsal and distillation, using both ground truth labels and logits stored in a memory buffer to maintain model performance across tasks. Evaluations on the Google Speech Command dataset show that DE-KWS outperforms existing CL baselines in average accuracy without increasing model size, offering an effective solution for resource-constrained edge devices. The scripts are available on GitHub  [^1] https://github.com/NeverlandCookies/DE-KWS for the future research.
@@ I Introduction
To address performance degradation, prior research has employed few-shot fine-tuning [11, 12, 13] to adapt KWS models to new scenarios using target-domain training data. However, this approach often leads to poor performance on the original source-domain data, a phenomenon known as catastrophic forgetting [14]. This occurs because new knowledge gradually replaces previously learned information.
To solve catastrophic forgetting, continual learning (CL) aims to continuously acquire new knowledge while retaining and reusing previously learned information. CL has been widely applied in speech processing [15, 16], audio classification, and other fields. It can be divided into task-incremental and class-incremental learning, depending on the presence of a task ID. Recent strategies like progressive continual learning [17] apply task-incremental continual learning to KWS, enabling models to adapt to new data while retaining previous knowledge. However, despite these advancements [18, 19], two key limitations remain the reliance on task-ID as auxiliary information, which is not always available, and the growing storage requirements as tasks increase. In this paper, we focus on class-incremental learning, which is more practical as it does not rely on task IDs. In continual learning, class-incremental learning (CIL) integrates new classes into a model’s learning architecture sequentially. Although class-incremental learning has a higher potential for real-world applications, CIL in KWS remains a less-explored area. Recently, Xiao et al. [20, 21] proposed a diversity-based sampler integrated with the self-distillation approach for future rehearsal, and Yang et al. [22] introduced a replay-based dual-memory multi-modal framework for continual KWS tasks. However, these approaches primarily focus on preserving knowledge from previous tasks and often overlook the richer context within speech clips. Dark experience replay (DER) [23], which stores past logits to distill experiences over the entire training trajectory, has been suggested as a method, but its application in the KWS domain remains unexplored.
In this paper, we present a CIL approach for KWS called Dark Experience for Keyword Spotting (DE-KWS). it relies on dark knowledge [24, 25] for distilling past experiences, sampled over the entire training trajectory. DE-KWS combines rehearsal and distillation, using both ground truth labels and dark knowledge embedded in the logits. We create a memory buffer that stores samples, their labels, and logits throughout the training process. The ground truth labels help the model remember past samples through experience replay, while the logits ensure the model retains its response to previous data. We tested DE-KWS on the Google Speech Command dataset [26], following prior work’s setup, and found that it outperforms other CIL baselines in Average Accuracy without increasing the network’s size. Our code is made available at the GitHub for future research.
@@ IV-B Ablation Study
IV-C Exploration of task numbers

### v2

@@ Abstract
Spoken keyword spotting (KWS) is crucial for identifying keywords within audio inputs and is widely used in applications like Apple Siri and Google Home, particularly on edge devices. Current deep learning-based KWS systems, which are typically trained on a limited set of keywords, can suffer from performance degradation when encountering new domains, a challenge often addressed through few-shot fine-tuning. However, this adaptation frequently leads to catastrophic forgetting, where the model’s performance on original data deteriorates. Progressive continual learning (CL) strategies have been proposed to overcome this, but they face limitations such as the need for task-ID information and increased storage, making them less practical for lightweight devices. To address these challenges, we introduce Dark Experience for Keyword Spotting (DE-KWS), a novel CL approach that leverages dark knowledge to distill past experiences throughout the training process. DE-KWS combines rehearsal and distillation, using both ground truth labels and logits stored in a memory buffer to maintain model performance across tasks. Evaluations on the Google Speech Command dataset show that DE-KWS outperforms existing CL baselines in average accuracy without increasing model size, offering an effective solution for resource-constrained edge devices. The scripts are available on GitHub  [^1] https://github.com/NeverlandCookies/DE-KWS for future research.
@@ I Introduction
To address performance degradation, prior research has employed few-shot fine-tuning [11, 12, 13, 14] to adapt KWS models to new scenarios using target-domain training data. However, this approach often leads to poor performance on the original source-domain data, a phenomenon known as catastrophic forgetting [15]. This occurs because new knowledge gradually replaces previously learned information.
To solve catastrophic forgetting, continual learning (CL) [16, 17, 18] aims to continuously acquire new knowledge while retaining and reusing previously learned information. CL has been widely applied in speech processing [19, 20, 21], audio classification, and other fields. It can be divided into task-incremental and class-incremental learning [22], depending on the presence of a task ID. Recent strategies like progressive continual learning [23] apply task-incremental continual learning to KWS, enabling models to adapt to new data while retaining previous knowledge. However, despite these advancements [24], two key limitations remain the reliance on task-ID as auxiliary information, which is not always available, and the growing storage requirements as tasks increase. In this paper, we focus on class-incremental learning, which is more practical as it does not rely on task IDs. In continual learning, class-incremental learning (CIL) [25, 26, 27] integrates new classes into a model’s learning architecture sequentially. Although it has a higher potential for real-world applications, CIL in KWS remains a less-explored area. Recently, Xiao et al. [28] proposed a diversity-based sampler integrated with the self-distillation approach for future rehearsal, and Yang et al. [29] introduced a replay-based dual-memory multi-modal framework for continual KWS tasks. However, these approaches primarily focus on preserving knowledge from previous tasks and often overlook the richer context within speech clips. Dark experience replay (DER) [30], which stores past logits to distill experiences over the entire training trajectory, has been suggested as a method, but its application in the KWS domain remains unexplored.
In this paper, we present a CIL approach for KWS called Dark Experience for Keyword Spotting (DE-KWS). it relies on dark knowledge [31, 32] for distilling past experiences, sampled over the entire training trajectory. DE-KWS combines rehearsal and distillation, using both ground truth labels and dark knowledge embedded in the logits. We create a memory buffer that stores samples, labels, and logits throughout the training process. The ground truth labels help the model remember past samples through experience replay, while the logits ensure the model retains its response to previous data. Following prior work’s setup, we tested DE-KWS on the Google Speech Command dataset and found that it outperforms other CIL baselines in average accuracy without increasing the network’s parameters. Our code is made available at the GitHub.
@@ IV-B Ablation Study
IV-C Exploration of increasing task numbers